OpenStudy (anonymous):

May I have some help solving inequalities? I will Medal for some assistance with these problems? 3/5=x/5 3+3/4=n/3 I have four more but that's a start.

OpenStudy (anonymous):

wait im gonna solve it

OpenStudy (anonymous):

I got the first one, the second and so on are far more complex.

OpenStudy (anonymous):

what are they?

OpenStudy (anonymous):

\[\frac{ 1 }{ 9y-7 } = \frac{ 2 }{ 4y }\]

OpenStudy (anonymous):

\[\frac{ 2 }{ x+5 }=\frac{ 1 }{ x+3 }\]

OpenStudy (trojanpoem):

18 y - 14= 4 y 14 y = 14 y = 1

OpenStudy (trojanpoem):

x+ 5 = 2x + 6 x = -1

OpenStudy (trojanpoem):

Buddy, to solve such problems just hit both side by the bottom of the fraction

OpenStudy (trojanpoem):

sides*

OpenStudy (trojanpoem):

this scheme will help you solve any problem of this type

OpenStudy (anonymous):

Right on, thanks for the advice.

OpenStudy (trojanpoem):

You are welcome.But if you stuck with a problem don't hesitate to ask for help

OpenStudy (anonymous):

I'll message you if need be, my assignment is only six problems long. The last one is confusing the heck out of me though.

OpenStudy (anonymous):

I'll just post it now and ask. It asks to solve for y... \[\frac{ x }{ y}=\frac{ a }{ b }\]

OpenStudy (trojanpoem):

ay = b x y = bx/a

OpenStudy (trojanpoem):

In such problem , don't expect to get numbers even if you were a wizard.

OpenStudy (anonymous):

Huh, okay right on, thanks man.

OpenStudy (trojanpoem):

yw